OpenStudy (sasogeek):

if 2+2=5 , 3+3=7 and 5+7=1 , what is 2+3 ? if 2+2=5 , 3+3=7 and 5+7=1 , what is 2+3 ? @Mathematics

OpenStudy (mertsj):

5+7 = 2+2+3+3 by substitution. So 2+2+3+3 = 1 since 5+7 = 1 So 2+3 = 1-2-3

OpenStudy (sasogeek):

answer is 2+3=0.5 since (2+3)+(2+3)=1

OpenStudy (anonymous):

2+2+3+3 = 1 2+3 = 1/2

OpenStudy (mertsj):

I'm not so sure, moneybird. This are obviously not the numbers we are familiar with or 2+2 WOULD NOT be 5 and 3+3 would not be 7. So how do we know there is such a think as division in this set of numbers? And furthermore, is there is, how do we know what 1/2 means?

OpenStudy (anonymous):

2+2=5 3+3=7 5+7 = (2+2)+(3+3) = 1 = 2(2+3) = 1 2+3 = 0.5

OpenStudy (anonymous):

I know, but the division rules and addition rules doesn't change

OpenStudy (mertsj):

Its not that I don't understand where you got it. How do you know those properties apply in this weird number set?

OpenStudy (anonymous):

Well, i am not for sure. But It's still math and this question is obviously testing the skills of your division and addition properties on algebra

OpenStudy (mertsj):

It could be testing your ability to absorb new unfamiliar definitions and apply them correctly. In other words it might be a logic problem and not an algebra problem.